OpenStudy (anonymous):

What is the solution of the equation? (x - 2)^3/4 = 8

OpenStudy (anonymous):

\((x-2)^\frac{3}{4}=8\\\left((x-2)^\frac{3}{4}\right)^\frac{4}{3}=8^\frac{4}{3}=(\sqrt[3]{8})^4=16\)

OpenStudy (anonymous):

\(x-2=16\\x=18\)